# REAR: A Relevance-Aware Retrieval-Augmented Framework for Open-Domain Question Answering

### v1

@@ 5.1 Experimental Setup
Datasets. We extensively test on four open-domain QA datasets: Natural Questions Kwiatkowski et al. (2019), TriviaQA Joshi et al. (2017), SQuAD Rajpurkar et al. (2016), and WebQuestions Berant et al. (2013). We follow the training/validation/test split in prior work Karpukhin et al. (2020) and report the results on the test set.
LLMs Natural Questions TriviaQA WebQuestions SQuAD
JAcc Hit EM F1 JAcc Hit EM F1 JAcc Hit EM F1 JAcc Hit EM F1
@@
[…]
Judge-then-Generate (4-shot)
[…]
Rank-then-Generate (4-shot)
[…]
@@
[…]
Table 1: A comparison between REAR and baselines on NQ, TriviaQA, WQ and SQuAD datasets. JAcc, Hit, EM and F1 are short for judgment accuracy, Hit@1 (Recall@1), exact match scores and F1 scores. The best and second-best results are in bold and underlined fonts respectively. ${}^{\dagger}$ is evaluated using accuracy (Acc) instead of EM, which is a less strict metric that measures whether the responses contain the answers. The last two lines are our REAR with different routing strategies: “w/ pr” means the path-reliability strategy, and “w/ kc” means the knowledge-consistency strategy.
@@
• Judge-then-generate: We following the approach of the prompting strategy Yoran et al. (2023). The LLMs are required to determine document relevance before generating responses. It employs 4-shot demonstrations (Fig. 5), and provides the top 1 retrieved document.
• Rank-then-generate: We first task the LLMs with ranking the documents they are most inclined to deem “relevant” from the top 10 retrieved documents, by assessing the perplexity of their relevance judgments. Subsequently, the document selected is utilized by the LLMs for RAG.
For open-source LLMs, we consider Llama2-Chat Touvron et al. (2023), Mistral-Chat Jiang et al. (2023), Baichuan2-Chat Yang et al. (2023), and ChatGLM3 Du et al. (2022).
(2) Specially designed RAG methods: we also consider RobustLM Yoran et al. (2023) and Self-RAG Asai et al. (2023) as baselines, which have been specially optimized for the RAG tasks. To ensure a fair comparison, we fine-tune these two models with our training set. Furthermore, the two frameworks above are evaluated with the same set of retrieved documents as used for REAR.
Metrics. We employ four metrics to evaluate the model’s capability of relevance assessment and QA accuracy. JAcc, which denotes judgment accuracy, measures the model’s accuracy in assessing if the documents are relevant. Consistent with prior research Khattab and Zaharia (2020), documents are labeled as “relevant” if they contain the answers. Hit is the short form of “Hit@1”, which evaluates if the document referenced for the model’s final answer generation is relevant. Exact Match (EMLee et al. (2019) calculates whether responses exactly match the gold truth answers. F1 calculates the precision-recall overlap of predicted and true answers.
@@ 5.2 Main Results
First, our REAR approach surpasses all the other baselines in coarse-grained relevance binary discrimination (higher JAcc), and fine-grained document relevance comparison (higher Hit), achieving the best generation performance (higher EM and F1 scores) across the four datasets.
Second, REAR achieves superior results than RobustLM which employs the same training data as ours. This highlights the strengths of our architecture and training method. Furthermore, both RobustLM and REAR outperform Self-RAG, indicating that our data construction strategy is a crucial factor in enhancing RAG performance.
Third, while LLM-based baselines demonstrate a natural proficiency in discerning document relevance under few-shot scenarios (as seen in the JAcc metrics for Judge-then-Generate and Rank-then-Generate), they fall short in the nuanced comparison of relevance (noted in the Hit metrics for Rank-then-Generate). This limitation hampers their ability to effectively utilize multiple documents, highlighting an important scenario where our REAR architecture and training approaches excel.
Finally, LLMs like ChatGLM3 and Mistral exhibit better performance based on individual documents rather than given the top-10 documents as a reference on Natural Questions. This suggests that filtering irrelevant retrieved documents could be another viable option for enhancing RAG, rather than simply increasing document quantity.
@@ 5.3 Detailed Analysis
In this part, we further present the analysis of the ablation study and the impact of retrieved documents.
@@
w/   Path-Reliability Inf. 67.48 52.91 61.49
w/o Bi-granularity Train. 66.54 51.88 59.91
w/o Noise-resistant Train. 49.25 25.54 33.05
w/o Rank Head Arch. 13.80 38.14 47.44
Table 2: Ablation study on our REAR. The “aspect” denotes the affected aspect. “Inf.”, “Train” and “Arch.” denotes inference, training and architecture respectively
5.3.1 Ablation Study
We analyze how each of the proposed components affects final performance. Table  2 shows the performance of our default method and its four variants in three aspects, including the architecture, training algorithm and inference strategy.
(1) w/o Rank Head: the variant without the integration of the ranking module. We utilize language generation to assess relevance degrees instead of the rank head. The document is selected based on the probability of generating judgmental statements. There is a notable drop in the comparison accuracy (see Hit@1 metric), similar shortfall is also observed in Self-RAG (Table 1). This demonstrates the effectiveness of our architectural design, which not only minimizes interference between language generation and relevance discrimination, but also facilitates the incorporation of various loss functions.
@@ 5.3.1 Ablation Study
(4) w/ Path-Reliability: using the path-reliability strategy instead of the knowledge consistency strategy. The path-reliability approach achieves higher Hit@1 rates, yet falls behind in EM and F1 scores compared to the knowledge-consistency strategy. The latter conducts a self-verification of outputs based on its generation ability, effectively integrating inherent knowledge in relevance assessment, which enhances the accuracy of RAG.
@@ 5.3.2 Impact of Retrieved Documents
[…]
@@ 6 Conclusion
In this paper, we aimed to enhance the self-awareness of source relevance in RAG systems, and proposed REAR, a RElevance-Aware Retrieval-augmented approach for open-domain question answering (QA). For model architecture, we explicitly incorporated a specially designed rank head to precisely capture the relevance signals, and employed it to guide the utilization of external knowledge. For model training, we designed an improved training method with bi-granularity relevance fusion and noise-resistant training, which enhance the capacities of fine-grained relevance assessment and adaptive use of retrieved documents. Extensive experiments on four datasets demonstrate the effectiveness of REAR’s relevance assessment and knowledge utilization.
As future work, we will extend the proposed approach REAR to dealing with more fine-grained source utilization (e.g., passage or sentence level augmentation), and also consider applying REAR to other knowledge-intensive tasks.
@@
Appendix A Illustration of Fine-Gained Relevance Comparison and Details in Labeling
@@
Given a query (Q) and the 5 retrieved documents (D1~D5) as shown in Fig. 3, traditional annotation methods in previous works typically use the presence of an answer within a document as the criterion to label it as positive or negative. In this illustration, both D1 and D2 meet this criterion. However, it’s observable that while D1 directly answers Q, D2 merely states that Wilhelm was awarded for discovering X-rays, requiring some external knowledge for inference. Training models on simple binary classification does not differentiate the superiority of D1 over D2, potentially leading to inaccuracies in finer relevance judgment.
Inspired by previous work RankGPT Sun et al. (2023) which attempts to train the rerankers by distilling GPT-4’s ranking results. We opt for a more economical solution by annotating with small-scale, well-trained cross-encoder rerankers. Nonetheless, these models may still harbor the potential for misjudgment. To address this, we linearly combine the binary label with cross-encoder scores, ensuring that documents with answers are prioritized over those without, and within the same binary category, documents with higher cross-encoder scores are ranked above those with lower scores. To mitigate noise from rerankers, we disregard differences smaller than 0.1 in fine-gained relevance training.
@@
We utilize four open-domain QA datasets, Natural Questions Kwiatkowski et al. (2019), TriviaQA Joshi et al. (2017), WebQuestions (WQ) Berant et al. (2013) and SQuAD Rajpurkar et al. (2016)).
@@
Training Data Combination. We followed the same train, test, and dev set splitting approach as DPR Karpukhin et al. (2020), In our study, we tailored the training dataset composition to optimize model performance across various question-answering tasks. Specifically, we integrated the complete datasets from Natural Questions (NQ), TriviaQA, and WebQuestions (WQ), applying a five-fold repetition to the WQ dataset to address its relatively smaller volume. For the SQuAD dataset, we opted to utilize only 2/9 of its training data, based on preliminary observations of its quicker convergence rates. Our detailed analysis (Section 5.3) explores the implications of employing a more straightforward strategy—merging the full training sets without repetition or selective sampling. Although this method simplifies dataset preparation, it yielded slightly lower performance compared to the aforementioned tailored approach.
Training Data Construction. We utilize task-specific dense retrievers to gather the top 100 retrieved documents from each dataset, creating a candidate document pool. Relevance scores for these documents are annotated using the reranker described in Appendix A. To avoid introducing false-positive and false-negative samples, we follow a strategy akin to RocketQA Qu et al. (2021), excluding positive samples with scores below 0.1 and negative samples with scores above 0.9. The dataset is constructed using the two highest-scoring positive documents and six negative documents sampled based on Eq. 16, with parameters $a$, $b$, and $k$ set to 15, 2, and 3, respectively.
Appendix C Details on Retrievers and Rerankers
Retrievers. For both training and inference, we employ advanced, task-specific retrievers to gather the necessary data. During the training phase, the RocketQAv2 Ren et al. (2021b) retriever is utilized for the NQ dataset, and a dense retriever as described by  Izacard and Grave (2021a) is used for the TriviaQA dataset. For the SQuAD and WQ datasets, we implement a strategy incorporating in-batch negatives and joint retriever-ranker training, starting from the Contriever-msmarco Izacard et al. (2021) checkpoint. In the inference stage, to ensure a relatively fair comparison with FiD Izacard and Grave (2021b), we adopt the dense retriever method outlined by  Izacard and Grave (2021a) to compile the top 10 documents. The retrievers we have trained are specifically applied during inference on the SQuAD and WQ datasets. The metrics of the retrieved documents for inference are shown in Table. 4.
@@ Appendix C Details on Retrievers and Rerankers
Rerankers. For the NQ dataset, we utilize the RocketQAv2 reranker due to its publicly available checkpoint. Electra-base Clark et al. (2019) serves as the backbone for training rerankers on the remaining three datasets, which are then employed for annotating, filtering, and sampling the training data.
Appendix D Training and Inference Details
Training Settings Following the methodology of Self-RAG Asai et al. (2023), our training employs a two-stage process. Initially, we focus on optimizing the LLM with coarse-grained loss as defined in Eq. 10. In the next stage, we apply joint optimization combining relevance assessment and generation adjustment, as specified in Eq. 14. The training utilizes a learning rate of 1e-6, a warm-up ratio of 0.03, a batch size of 64 and a cosine scheduler across two epochs. Our experiments leverage the computational power of 8 NVIDIA Tesla A100 GPUs, each with 40G of memory.
Inference settings. During the inference phase, we configure the threshold $\gamma$ to 13 for binary discrimination of document relevance in Eq. 8 , and set $\lambda$ to 9 to act as the knowledge-consistency scaler in Eq. 9. The generation process employs greedy generation.
Knowledge: {retrieved document 1} {retrieved document 2} …… {retrieved document 9} Answer the following question with a very short phrase, such as “1998”, “May 16th, 1931”, or “James Bond”, to meet the criteria of exact match datasets. Question: {question} Answer:
Figure 4: Instruction format for “direct RAG QA”.
@@ Appendix D Training and Inference Details
Figure 5: Instruction format for “judge-then-generate” and “rank-then-genrate”.
Appendix E Details on Baselines
All open-source models employed in this study are obtained from huggingface, with each model’s respective chat instruction format used for generation. No system instructions are utilized in any of the tests, and the greedy decoding strategy is applied for generation. The specific instruction formats used in our tests are illustrated in Fig. 4 and Fig. 5. In the “rank-then-generate” setup, we calculate the negative log perplexity $(-\log(ppl))$ for generating “yes” and “no”, ranking them based on this metric to select the model that maximizes the probability of generating “yes” over “no”.
Appendix F Computational Efficiency of Training and Inference
In this section, we discuss the computational efficiency of our model during both the training and inference phases. Our model introduces new modules and splits the inference process into two steps: relevance discrimination and answer generation. We employed various efficient fine-tuning and inference techniques, ensuring our model’s computational demands during training and inference are comparable to those of the original LLM architecture.
Training Process. During training, we introduced negative examples in both bi-granularity fusion and noise-resistant training (Section  4.2), allowing the same instruction data to optimize two objectives simultaneously. This approach does not significantly increase computational overhead, akin to introducing a new word into the vocabulary to LLMs. Additionally, we adapted flash-attention Dao et al. (2022); Dao (2023) techniques to accelerate training.
Inference Process. In the inference phase, we utilized the key-value (KV) cache Vaswani et al. (2017) technology, making the computational cost of relevance discrimination equivalent to generating an extra token in the generation process. We also adapted vLLM Kwon et al. (2023), leveraging PagedAttention for efficient inference.

### v2

@@ 5.1 Experimental Setup
Datasets. We collect the training data from the Natural Questions (NQ) Kwiatkowski et al. (2019) training set. To ensure the model’s adaptability, we also test its performance on three additional open-domain datasets, including TriviaQA Joshi et al. (2017), WebQuestions (WebQ) Berant et al. (2013), and SQuAD Rajpurkar et al. (2016), showing its generalization capabilities to out-of-domain data. We follow the test split in prior work Karpukhin et al. (2020). The details are in Appendix B.
LLMs NQ TriviaQA WebQ SQuAD Average
EM F1 EM F1 EM F1 EM F1 EM F1
@@
RobustLM prompting (4-shot)
@@
Table 3: A comparison between REAR and baselines on NQ, TriviaQA, WebQ and SQuAD datasets and the averaged performance. our REAR approach surpasses all the other baselines in QA performance. The best and second-best results are in bold and underlined fonts respectively. Self-RAG is evaluated using accuracy (Acc) instead of EM, which is a less strict metric that measures whether the responses contain the answers. The last two lines are our REAR with different verification strategies: “w/ Source Rel.” means the source-reliability strategy, and “w/ Knowledge Con.” means the knowledge-consistency strategy.
@@
• RobustLM prompting: We following the approach of the prompting strategy Yoran et al. (2023). The LLMs are required to determine document relevance before generating responses. It employs 4-shot demonstrations (Fig. 6), and provides the top 1 retrieved document.
For open-source LLMs, we consider LLaMA2-Chat Touvron et al. (2023), Mistral-It Jiang et al. (2023), Baichuan2-Chat Yang et al. (2023), and ChatGLM3 Du et al. (2022).
(2) Specially designed RAG methods: we also consider fine-tuned RobustLM Yoran et al. (2023) and Self-RAG Asai et al. (2023) as baselines, which have been specially optimized for the RAG tasks. To ensure a fair comparison, the two frameworks above are evaluated with the same set of retrieved documents as used for REAR.
Metrics. We employ three metrics to evaluate the model’s capability of QA accuracy. Exact match (EMLee et al. (2019) and F1 are widely adopted for open-domain QA evaluation. EM calculates whether responses exactly match the gold truth answers and calculates the precision-recall overlap of predicted and true answers. We further evaluate the accuracy in determining the relevance of the given document for LLMs with another two metrics. Hit@1 evaluates if the document referenced for the model’s final answer generation is relevant. JAcc, short for judgmental accuracy, measures the proportion of documents correctly evaluated by LLMs as relevant or not.
@@
[…]
JAcc Hit@1 JAcc Hit@1 JAcc Hit@1 JAcc Hit@1
+ RobustLM-prompting Yoran et al. (2023) 25.04 - 43.36 - 28.84 - 16.84 -
+ RobustLM-training Yoran et al. (2023) 56.59 - 56.09 - 49.61 - 56.99 -
+ Self-RAG Asai et al. (2023) 19.81 51.11 35.69 64.47 25.69 47.98 10.73 38.73
+ REAR (ours) 74.04 66.79 80.79 74.98 65.99 56.69 59.36 53.26
Table 4: The relevance discrimination and comparison capabilities of REAR and previous approaches. Generative LLMs struggle to determine the relevance degree of the given document, while our REAR overcomes it with the well-designed assessment module.
@@ 5.2 Main Results
First, our REAR approach surpasses all the other baselines in QA performance. REAR not only performs well on the trained dataset, but also achieves good results on non-training datasets. This demonstrates that our precise signals for capturing relevance effectively guide the generation process. Thus, LLM can generate with good use of both parametric and external knowledge.
Besides, the result shows the efficiency of data construction method, even without access to GPT APIs. Self-RAG labels the relevance degree with GPT-4, while RobustLM and REAR utilize our proposed label machine and sampling method. The result indicates that our data construction strategy is effective and less costly.
Third, generative LLMs struggle to determine the reliability degree of the given document, while our REAR overcomes it with the well-designed assessment module. As shown in Table 4, even the fine-tuned generative approaches (RobustLM-training and Self-RAG) adequately discriminate relevance. In comparison, REAR significantly enhances this capability, highlighting its effectiveness in architectural design.
@@ 5.3 Detailed Analysis
In this part, we further present the analysis of the ablation study and impacts of retrieved documents.
5.3.1 Ablation Study
We analyze how each of the proposed components affects final performance. Table  5 shows the performance of our default method and its five variants in three aspects, including the architecture, training objectives and sampling strategy.
(1) w/o Assessment: the variant without the integration of the rating module. We utilize language generation to assess relevance degrees instead of the rank head. The document is selected based on the probability of generating judgmental statements. There is a notable drop in the comparison accuracy (see Hit@1 metric), similar shortfall is also observed in Self-RAG (Table 3). This demonstrates the effectiveness of our architectural design, which not only minimizes interference between language generation and relevance discrimination, but also facilitates the incorporation of various loss functions.
(2) w/o Consistency: using the path-reliability strategy instead of the knowledge consistency strategy. The path-reliability approach achieves higher Hit@1 rates, yet falls behind in EM and F1 scores compared to the knowledge-consistency strategy. The latter conducts a self-verification of outputs based on its generation ability, effectively integrating inherent knowledge in relevance assessment, which enhances the accuracy of RAG.
@@ 5.3.1 Ablation Study
w/o Assessment Arch. 13.80 38.14 47.44
w/o Consistency Arch. 67.48 52.91 61.49
w/o Bi-granularity Obj. 66.54 51.88 59.91
w/o Noise-resistant Obj. 49.25 25.54 33.05
w/o Sampling Sam. 61.99 49.00 53.62
Table 5: Ablation study on our REAR. The “aspect” denotes the affected aspect. Arch., Obj. and Sam. denote architecture, training objective and sampling strategy.
@@
(5) w/o Sampling: the variant training with random hard negatives for training. We can observe a significant drop in relevance assessment capability, further illustrating the effectiveness of our method.
@@ 5.3.2 Impact of Retrieved Documents
Figure 3: Results of RAG performance vary in overall document count and quality. The left one presents RAG performance with varying numbers of retrieved documents. The right one is the results of RAG with different retriever engines. R1, R2, and R3 represent BM25, Contriever-msmarco, and the FiD-distilled retriever, R1 < R2 < R3 (Table 9 of the Appendix) .
@@ 6 Conclusion
In this paper, we aimed to enhance the self-awareness of source relevance in RAG systems, and proposed REAR, a RElevance-Aware Retrieval-augmented approach for open-domain question answering (QA). For model architecture, we explicitly integrate an assessment module to precisely capture the relevance signals, and employ it to guide the utilization of external knowledge. For model training, we designed an improved training method with bi-granularity relevance fusion and noise-resistant training, which enhance the capacities of fine-grained relevance assessment and adaptive use of retrieved documents. Our data construction strategy collects high-quality data without access to GPT APIs. Extensive experiments on four datasets demonstrate the effectiveness and generalization of REAR’s knowledge utilization.
As future work, we will extend the proposed approach REAR to deal with more fine-grained source utilization (e.g., passage or sentence level augmentation), and also consider applying REAR to other knowledge-intensive tasks.
@@
Appendix A Details on Fine-Gained Relevance Optimization
@@
We first illustrate why to design the fine-grained optimization for the assessment module. Traditional annotation methods always use a binary labeling method Karpukhin et al. (2020), which is based on the presence of an answer within a document. As shown in Fig. 4, both D1 and D2 are labeled as “relevant”. However, while D1 allows for direct answer derivation, D2 requires additional external knowledge for induction. Training models solely on simple binary classification fails to distinguish the superiority of D1 over D2, potentially leading to inaccuracies in finer relevance judgments.
Previous work has achieved success in relevance assessment by distilling the ranking results from GPT-4 Sun et al. (2023). Inspired by it, we propose a less costly solution by labeling with small-scale, well-trained cross-encoder rerankers RocketQAv2 Ren et al. (2021b). Despite the good relevance evaluation performance, it still may get wrong. We adopt three strategies to reduce the negative impact of annotation errors on training. Firstly, we design the sampling method (Eq. 11), which reduces the likelihood of potentially false negatives being sampled. Besides, we linearly combine the binary label with cross-encoder scores. Thirdly, to mitigate noise from rerankers, we disregard differences smaller than 0.1 in fine-gained relevance training in Eq. 7. These strategies enhance the quality of training data, which in turn improves the performance of REAR.
@@
We utilize four open-domain QA datasets, Natural Questions (NQ) Kwiatkowski et al. (2019), TriviaQA Joshi et al. (2017), WebQuestions (WebQ) Berant et al. (2013) and SQuAD Rajpurkar et al. (2016)).
Dataset NQ TriviaQA WebQ SQuAD
Num. Data 3,610 11,313 2,032 10,570
Table 7: Dataset statistics of the test set.
@@
NQ is used for both training and inference, while the other three are only used for inference. We use the same split as previous work Karpukhin et al. (2020). The training set of NQ contains 58,880 samples.
Appendix C Details on Document Collection
In this part, we introduce the retrievers we used to collect documents. We employ task-specific retrievers to acquire the retrieved document. For inference, we utilize FiD-distilled retrievers Izacard and Grave (2021a) for NQ and TriviaQA datasets. And we implement a strategy incorporating in-batch negatives and joint retriever-ranker training, starting from the Contriever-msmarco Izacard et al. (2021) checkpoint for SQuAD and WQ datasets. The recall and MRR rates of the retrieved documents for inference are shown in Table. 8.
@@ Appendix C Details on Document Collection
Appendix D Details on Implementation
Following the previous work Asai et al. (2023), we apply joint optimization combining relevance assessment and relevance-guided generation, as specified in Eq. 12. The training utilizes a learning rate of 1e-6, a warm-up ratio of 0.03, a batch size of 64 and a cosine scheduler for 1 epoch. Our experiments leverage the computational power of 8 NVIDIA Tesla A100 GPUs, each with 40G of memory.
Knowledge:
{retrieved document 1}
{retrieved document 2} …… {retrieved document 9} Answer the following question with a very short phrase, such as “1998”, “May 16th, 1931”, or “James Bond”, to meet the criteria of exact match datasets. Question: {question} Answer:
Figure 5: Prompts for “direct RAG QA”.
Appendix E Details on Baselines
In this part, we detail the prompt design and inference details for baselines. For the prompt-based inference, we utilize the instruction-tuned open-source models obtained from Hugging Face. Following the existing work Ren et al. (2024); Tang et al. (2024); Asai et al. (2023), we use the greedy decoding strategy for inference. The specific instruction formats used in our tests are illustrated in Fig. 5 and Fig. 6.
@@ Appendix E Details on Baselines
Figure 6: Prompts for “RobustLM based prompting”.